OpenStudy (anonymous):

15,400=1200(1+.04)^t

OpenStudy (anonymous):

so you want to find t?

OpenStudy (anonymous):

(1 + 0.04)^t = 15400/1200 t (ln (1.04) = ln (15400/1200) t =ln (15400/1200) / ln (1.04 t = 65.07